In a high-temperature reaction, hydrogen bromide and oxygen combine to produce water and bromine.
4HBr(g) + O2(g) ⭢ 2H2O(g) + 2Br2(g)
The rate expression was determined by experiment to be
Rate = k[HBr] [O2]
In one experiment in which the concentrations of both reactants were 0.250mol L-1, the rate was found to be 6.54 X 10-3mol L-1 s-1
Calculate the value of k with appropriate units.
1
Expert's answer
2020-11-02T09:03:35-0500

[Hbr] = 0.25 M

[O2] = 0.25 M

Rate = k[HBr][O2]

"6.54 \times 10^{-3} \;mol/L\;s = k(0.25 \;mol/L)(0.25 \;mol/L)"

"k = \frac{6.54 \times 10^{-3} \;mol/L\;s}{(0.25 \;mol/L)(0.25 \;mol/L)}"

"k = 0.104 \;L/mol\;s"